A quantity of 1.922 g of methanol (CH3OH) was burned in a constant-volume bomb calorimeter. Consequently, the temperature rose by 5.52°C. If the heat capacity of the bomb plus water was 8.75 kJ / °C, calculate the molar heat of combustion of methanol.
Answer #1

Heat = C×(∆T)

q = 8.75×5.52

= 48.3 J

∆U = -q/n

= -48.3×1000×32/1.992

= 776 kJ/mol